HUBBLE SPACE TELESCOPE DAILY REPORT #1493 PERIOD COVERED: 0000Z (UTC) 10/10/95 - 0000Z (UTC) 10/11/95 Daily Status Report as of 284/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 6119 (Snapshot Survey for Companions = of Planetary Nebula Nuclei) The WF/PC-2 was used to make snapshot survey observations of = planetary nebulae IC2448 and NGC-6210 to study their nuclei and to look = for possible companions. The observations were executed as scheduled, and = no anomalies were noted. 1.2 Completed FGS 6037 (Continuation of an Astrometric Search for = Planetary Companions to Proxima-Cent) The FGSs were used in the POS mode to make astrometric observations = of the star Proxima-Cent and a sub-set of eight reference stars. This is = to look for the presence of a planetary companion. The observations were = executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrent with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 6113 (Searching for Low Mass Stars: The Mass = Function ) The WF/PC-2 was used to make observations of globular cluster = NGC-6254 to search for very low mass (VLM) stars. The intent is to derive = the luminosity and mass functions for VLM stars. The observations were = completed as planned, and no problems were reported. 1.5 Completed WF/PC-2 6250 (WF/PC-2 Cycle 5 Calibration: Internal = Monitor) The WF/PC-2 was used to make a series of internal calibration = observations. Internal flats, biases, and K-Spot calibration images were = taken. These observations are designed to monitor the health of the = cameras. The observations were executed as scheduled, and no anomalies = were noted. 1.6 Completed GHRS/1 6070 (Interstellar Tin: Probing Galactic S-Process = Element Enrichment) The High Resolution Spectrograph (Side One) was used to make = Echelle A observations of star HD37128 in order to probe for interstellar = tin and to study the galactic S-process neutron capture of element = enrichment. The observations were completed as planned, and no problems = were reported, although the acquisition at 283/145113Z, just prior to this = proposal, failed to fine lock on FGS-1 only. HSTAR 5329 documented this = event. 1.7 Completed WF/PC-2 5949 (High Resolution Imaging of Galaxies = Responsible for QSO Lyman Alpha Absorption.) The WF/PC-2 was used to make high resolution observations of Field = 0531 in order to study galaxies which are responsible for intermediate = redshift QSO Lyman Alpha absorption. It is the extended gaseous regions = around these galaxies that produce the absorption. The observations were = executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 5998 (The Evolution of Massive Stars: Closing the = Loop Observationally) The WF/PC-2 was used to make observations of (OB Class) massive = stars in M-33 Field 7 which is located in the Andromeda galaxy. The goal = is to study the evolution of this class of massive stars. The observations = were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 5999 (The Nuclear Structure of S0 Galaxies: A PC = Imaging Survey) The WF/PC-2 was used to make observations of S0 class galaxy = NGC-1389 in order to study its nuclear structure. This is a part of a = large survey of S0, E/S0, and S0/a class galaxies. The targets will cover = a wide range in luminosity and environment. The observations were = completed as planned, and no anomalies were noted. 1.10 Completed GHRS/2 6065 (ZAMS Atmospheric Structure as a Function of = Rotation) The High Resolution Spectrograph (Side Two) was used to make = emission line spectrographic observations of the zero-age main sequence = (ZAMS) star HD82558 in order to study its atmospheric structure as a = function of rotation. Wavelength calibrations, internal flats, and dark = calibration images were also taken. The observations were executed as = scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 The FGS acquisition at 283/145113Z failed to fine lock back-up on FGS-1 only. This fact was recorded in HSTAR 5329. Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 17 Successful : 17 2.3 Operations Notes: Operations personnel are in the process of closely monitoring the HGA gimbal temperatures during a period of high exposure. Peak temperatures as of 284/0935Z were approximately 44 degrees C for the X-axis and approximately 42 degrees C for the Y-axis. ESTR #1 was re-spooled at 284/0115Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-